OpenStudy (anonymous):

Using the equation you found for “double my age then increase it by 12, the result is 62” and the ages 23, 25 and 27, showing work for all three values, tell which is a solution and which is not a solution to the equation.

OpenStudy (love_ranaa):

Double my age then + 12 is like saying age x age then add 12 to that..So let's start with 23.

OpenStudy (love_ranaa):

23 x 2= 46 + 12 = 58

OpenStudy (love_ranaa):

25 x 2 = 50 +12 = 62

OpenStudy (love_ranaa):

27 x 2 = 54 + 12= 66

OpenStudy (anonymous):

Thank you :D